SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Friday 6 December 2002

Status Report From: Goddard Space Flight Center
Posted: Friday, December 6, 2002

ANNOUNCEMENTS
-------------
* Current RHESSI default target is AR 10212.

* There will be VDS (Voice Distributuion System) maintenance
  on Saturday, 7 December from 8 a.m. until 3 p.m. EST. No
  impact is expected.
 
* Reminder to instrument teams:  Please send updated contact
  lists to the SOCs.  Phone numbers have changed for some,
  and the holiday vacation dates should be noted as well.


FOT REPORT
---------- 
Spacecraft Status: Nominal

Spacecraft Anomalies: 
338/1734  UVCS entered "SAFE" mode, cause TBD.  UVCS was in 
          "SAFE" mode from 1734 until 2047.  Impacts TBD.

Accomplished Activities: 
DOY 338:        VIRGO, New tracking star, mag = 6.5,
                SUMER powered down internally, halted operations,
                Switch to DHSS Science Submode 6,
                Eclipse operations compelted successfully

DOY 339:        VIRGO, Nominal HGA Table, CAE-A in AOCS HW Config Word

Planned Activities: VIRGO, New control star, mag 6.5, New 
		    tracking star, mag = 5.5

Upcoming Operations: None

Ground Anomalies: 
338/1823        Incorrect TLM timestamp from 1823 to 1945 due to a 
                TCA failure at D46.
338/2012        NRT commaning delayed 9 minutes due to a CMS crash 
                in the POCC.
338/2254        SH4XT1 terminated TSTOL session, cause unknown.  
                Display terminal inoperative from 2254 to 2258, and 
                again from 2336 until 2337.
339/0725        Incorrect TLM timestamp from 0725 to 0732 due to a 
                TGC failure at D66.

Other: TDRS Launched successfully Wednesday night.  No schedule 
       impacts to SOHO for Week 49.


SOLAR STATUS
------------
12/04 DOY 338 - 12/06 DOY 340

 EIT daily On-Line Solar Status and observations:
   http://umbra.nascom.nasa.gov/eit/plan/log_form20021206.html

 LASCO daily On-Line Solar Status and observations:
   http://lasco-www.nrl.navy.mil/observations/Todays_obs.html

 Web Page for Planning:
   http://www.bbso.njit.edu/arm



INSTRUMENT STATUS
-----------------
CDS:    Nominal.  Particle beam study, line widths at the equatorial coronal
	hole. Weekend: usual studies, energetic event watch, CME onset 
	watch and off-limb spectral studies (North- Sat; South - Sun).

UVCS:	Nominal. Coronal hole and synoptic today through the weekend. 

LASCO:  Nominal. C2 and C3 synoptics, JOP159.

EIT:    Nominal. 195 CME watch, synoptics, JOP159.

MDI:    Nominal. Full disk cropped magnetograms and dopplergrams.

TRACE:  TRACE is out of safe hold and will observe AR 10207 when 
	observations resume at 16 UT today.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter